OpenStudy (anonymous):

Gina's age is 2 times more than 3 times Andrews age if the sum of their ages is 54 years, how old is Gina?

OpenStudy (anonymous):

g = 2 * 3 * a g + a = 54

OpenStudy (texaschic101):

g + a = 54 g = 2(3a) = 6a now sub 6a in for g in the first equation. This will give you a (Andrew's age). Then sub that into either of the equations to find g (Gina's age)